OpenStudy (anonymous):

A cylinder has a radius of 12 cm and a height of 10 cm. What is its volume? Select one: a. 240pi cm^3 b. 1200pi cm^3 c. 1440pi cm^3 d. 2400pi cm^3 I got 4000 and something so im lost

OpenStudy (mrnood):

what is the area of one end of the cylinder? (leave pi in the answer...)

OpenStudy (calculusxy):

\[\huge B \times h\] B means the area of the base.. in other words out formula is like: \[\huge (\pi r^2) \times h\]

OpenStudy (mrnood):

@Legendary maybe you have got the answer right - but if you look at the answers they have left pi in the answer have you maybe multiplied pi * r^2 * h (that would give 4000 something...)

OpenStudy (anonymous):

i got 4,523

OpenStudy (anonymous):

i dont get it

OpenStudy (mrnood):

read my post - your answer IS correct - but the answer given has pi left as pi - not multiplied out by 3.14...

OpenStudy (anonymous):

okay so what do i do

OpenStudy (calculusxy):

we are doing it in terms of pi. in other words, we will work with the radius and the height in the formula, but not the pi. r^2 = 12^2 = 144 height = 10 144 x 10 = 1,440 and we will attach pi to 1440 \[\large 1,440\pi\]

OpenStudy (anonymous):

thank you guys